### v1

# Magnetic Materials for Quantum Magnonics

Abstract
Quantum magnonics studies the quantum properties of magnons — the quanta of spin waves — and their application in quantum information processing. Progress in this field depends on identifying magnetic materials with characteristics tailored to the diverse requirements of magnonics and quantum magnonics. For single-magnon excitation, its control, hybrid coupling, and entanglement, the most critical property is the ability to support long magnon lifetimes. This perspective reviews established and emerging magnetic materials — including ferromagnetic metals, Heusler compounds, antiferromagnets, altermagnets, organic and 2D van der Waals magnets, hexaferrites, and in particular yttrium iron garnet (YIG) — highlighting their key characteristics. YIG remains the benchmark, with bulk crystals supporting sub-microsecond Kittel-mode lifetimes and ultra-pure spheres achieving $\sim$18 µs for dipolar-exchange magnons at millikelvin temperatures. However, thin YIG films on gadolinium gallium garnet (GGG) substrates suffer from severe lifetime reduction due to substrate-induced losses. In contrast, YIG films on a new lattice matched, diamagnetic alternative, yttrium scandium gallium/aluminum garnet (YSGAG), overcomes these limitations and preserves low magnetic damping down to millikelvin temperatures. These advances provide a practical pathway toward ultralong-living magnons in thin films, enabling scalable quantum magnonics with coherent transport, strong magnon-photon, magnon-qubit coupling, and integrated quantum networks.
preprint: AIP/123-QED
I Introduction into quantum magnonics
Quantum computing promises transformative advances in computation, with potential impacts on cryptography, materials science, and artificial intelligence Nielsen and Chuang (2010); Arute et al. (2019). A central challenge for solid-state quantum technology is to identify an on-chip platform with a suitable nanoscale information carrier Chumak et al. (2022). Magnons — the bosonic quanta of spin-wave excitations in ordered magnets — are promising for this role. The emerging field of quantum magnonics Lachance-Quirion et al. (2019); Li et al. (2020, 2022); Chumak et al. (2022); Yuan et al. (2022) operates with quantum states of magnons, including single magnons, typically at millikelvin temperatures where thermal magnons are suppressed. Magnons cover gigahertz to terahertz frequencies Levchenko et al. (2024), scale to nanometer footprints Wang et al. (2019); Heinz et al. (2020), interface naturally with microwave circuitry Chumak, Serga, and Hillebrands (2017), and exhibit strongly nonlinear and nonreciprocal dynamics Chumak, Serga, and Hillebrands (2014); Wang et al. (2024); Pirro et al. (2021); Wang et al. (2023a); Zenbaa et al. (2025). These properties position magnons as promising candidates for quantum information processing, coherent transport, and entangled-state preparation Lachance-Quirion et al. (2019); Li et al. (2020, 2022); Chumak et al. (2022); Yuan et al. (2022).
At room temperature (RT), macroscopic quantum phenomena in magnon gases are already accessible. Theory and experiment have established Bose–Einstein condensation of magnons Demokritov et al. (2006); Serga et al. (2014); Schneider et al. (2020, 2021), magnonic supercurrents Bozhko et al. (2016); Kreil et al. (2018), and the magnonic Aharonov–Casher effect Serha et al. (2023); Nakata et al. (2014). In antiferromagnets, quantum magnonics can reach the terahertz regime. Cavity coupled antiferromagnetic magnons have been demonstrated in collinear and canted phases Boventer et al. (2023), and terahertz cavity magnon polaritons have been realized in prototype platforms Kritzell et al. (2023). Long distance coherent antiferromagnetic spin-wave propagation at RT further supports device scale viability Wang et al. (2023b). These advances complement broader arguments for magnons as quantum excitations with frequency tunability and compatibility with hybrid architectures Yuan et al. (2022); Chumak et al. (2022); Li et al. (2020).
Single-magnon operations in ferromagnetically ordered media require suppression of thermal magnon populations. At a frequency of 4 GHz, which is typical for superconducting qubits, high enough that three-magnon splitting is forbidden, and lifetimes are longer than at higher frequencies where viscous damping is strongerGurevich and Melkov (1996). To keep the thermal occupation below 0.01 by following the Bose–Einstein statistics, experiments must operate in dilution refrigerators at temperatures of about 40 mK or lowerClerk et al. (2010). The objective is to generate, control, and read out single magnons in hybrid systems that couple to photons, microwave cavities, phonons, or superconducting qubits Tabuchi et al. (2015); Wolski et al. (2020); Lachance-Quirion et al. (2020); Baity et al. (2021); Zuo et al. (2024).
Table 1: Comparison of magnetic media relevant to quantum magnonics.
Property YIG Cherepanov, Kolokolov, and L’vov (1993); Serha et al. (2025a); Bertaut and Forrat (1956); Serga, Chumak, and Hillebrands (2010); Serha et al. (2025b); Geller and Gilleo (1957); Klingler et al. (2017, 2014); Maier-Flaig et al. (2017) Permalloy (Py)Demidov and Demokritov (2015); Kalarickal et al. (2006); Sebastian et al. (2015); Patton (1968) CoFeB Brunsch (1979); Liu et al. (2011); Conca et al. (2013); Zenbaa et al. (2025) Hematite Chen et al. (2025); Hamdi, Posva, and Grundler (2023); Wang et al. (2023b); Samuelsen and Shirane (1970) Hexaferrite BaM Song et al. (2010); Harris (2012); Malkinski (2012); Zhu et al. (2021); Popov et al. (2022) Heusler CMFSLiu et al. (2009); Trudel et al. (2010); Oogane et al. (2010); Sebastian et al. (2012) CrPS de Wal et al. (2023); Bud’ko et al. (2021); Lee et al. (2017); Freeman et al. (2025)
Chemical composition $\mathrm{Y_{3}Fe_{5}O_{12}}$ $\mathrm{Ni_{81}Fe_{19}}$ $\mathrm{Co_{40}Fe_{40}B_{20}}$ $\mathrm{Fe_{2}O_{3}}$ $\mathrm{BaFe_{12}O_{19}}$ $\mathrm{Co_{2}Mn_{0.6}Fe_{0.4}Si}$ $\mathrm{CrPS_{4}}$
Structure single-crystalline polycrystalline amorphous single-crystalline single-crystalline single-crystalline 2D van der Waals
Gilbert damping $\alpha$ $3\times 10^{-5}$ $7\times 10^{-3}$ $4\times 10^{-3}$ $5\times 10^{-5}$ $7\times 10^{-4}$ $3\times 10^{-3}$ $\times 10^{-2} 11$1Used in micromagnetic simulationsFreeman et al. (2025).
Saturation magnetization @ RT $M_{\mathrm{{}_{S}}}$ (kA/m) 140 800 1250 320-390 1000
Exchange constant @ RT $A$ (pJ/m) 3.6 16 15 102010-20 6-9 13 10  [^1] Used in micromagnetic simulationsFreeman et al. (2025).
Curie (Néel) temperature $T_{\mathrm{{}_{C(N)}}}$ (K) 560 550870550-870 1000 950 726 >985>985 37
Theory identifies roles for magnons as quantum interconnects and as probes of solid state quantum effects Kounalakis, Bauer, and Blanter (2022); Jiang et al. (2023); Dols et al. (2024); Dey et al. (2025). Further proposals predict steady state magnon squeezing and photon-magnon entanglement in cavity hybrids using Kerr nonlinearities and engineered couplings Yang et al. (2021); Toklikishvili et al. (2023); Sharma, Viola Kusminskiy, and Bittencourt (2024). Magnomechanical schemes show how squeezing can enhance correlations and light-matter control and enable nonclassical state preparation Lu et al. (2023); Zhang et al. (2024). Experimentally, key milestones include coherent coupling of fundamental magnon modes to superconducting resonators Lachance-Quirion et al. (2019); Li et al. (2022); Song et al. (2025a), qubit based quantum sensing Wolski et al. (2020), single magnon detection with a superconducting qubit Lachance-Quirion et al. (2020), full Wigner function tomography of a single magnon Xu et al. (2023), and NV based platforms that quantify magnon mediated NV-NV coupling Fukami et al. (2024); Borst et al. (2023). Although typical Kittel-mode magnon lifetimes Lachance-Quirion et al. (2020); Xu et al. (2023) of several hundred nanoseconds are shorter than the decoherence times of state-of-the-art superconducting qubits, which reach tens of microsecondsKrantz et al. (2019), they are sufficient for entanglement transfer and for interfacing heterogeneous quantum systems Tabuchi et al. (2015); Li et al. (2020); Yuan et al. (2022).
The most recent single-magnon experiments have been carried out on bulk spheres of yttrium iron garnet (YIG) Lachance-Quirion et al. (2020); Xu et al. (2023), which provides the longest coherence with Kittel-mode lifetimes generally below one microsecond Serga, Chumak, and Hillebrands (2010), underscoring its importance as the leading platform for quantum magnonics. Significant progress has been made in extending lifetimes at low temperatures, at which in ultra-pure bulk YIG, dipolar-exchange magnons reach about 18 µs at millikelvin temperatures Serha et al. (2025b).
Scaling from localized resonators to circuit-level links requires propagating magnons. Propagation enables spatial separation of single-magnon sources and detectors and allows the transport of quantum states across a chip Van Loo, Morris, and Karenowska (2018); Karenowska et al. (2015); Knauer et al. (2023). This shifts attention from bulk magnetic resonators to thin films, where dipolar-driven magnons can travel. However, thin films perform worse than bulk samples, as their higher surface-to-volume ratio amplifies the effect of surface defects. A further challenge in thin YIG films arises from the paramagnetism of GGG substrates at low temperatures Serha et al. (2025c, 2024); Guo et al. (2022); Jermain et al. (2017); Mihalceanu et al. (2018); Kosen et al. (2019). This limitation has recently been overcome by growing YIG films on a newly developed lattice-matched diamagnetic substrate Guguschev et al. (2025); Serha et al. (2025d). These advances enable on-chip platforms that excite and detect propagating magnons with planar resonators Wang et al. (2024). They also connect to ultralow-temperature studies of damping in bulk YIG Serha et al. (2025b); Kosen et al. (2019), to wavenumber-dependent damping in films Schmoll et al. (2025a), and to nanoscale propagation at cryogenic temperatures Knauer et al. (2023). Together, this progress establishes a foundation for quantum magnonic circuits that integrate superconducting qubits, photons, and phonons and exploit the nonlinear and nonreciprocal dynamics of magnons Barman et al. (2021); Flebus et al. (2024); Chumak et al. (2015, 2022).
II Available magnetic materials
Magnonics relies on ordered magnetic media that efficiently support spin waves. In this Perspective we briefly present and analyze different material platforms, including those that have been widely used for many years as well as more recent and highly attractive candidates. An overview is provided in Tab. 1, and representative examples are illustrated in Fig. 1.
Ferromagnetic metals
A common class of materials used to excite spin waves is ferromagnetic metals, valued for straightforward fabrication and high saturation magnetization (see Tab. 1). The most common examples are Permalloy ($\mathrm{Ni_{81}Fe_{19}}$)Demidov and Demokritov (2015); Kalarickal et al. (2006); Sebastian et al. (2015); Patton (1968), CoFeBTalmelli et al. (2020); Wojewoda et al. (2024); Brunsch (1979); Liu et al. (2011); Conca et al. (2013); Zenbaa et al. (2025) ($\mathrm{Co_{40}Fe_{40}B_{20}}$) and CoFeSchoen et al. (2016). However, their use in quantum magnonics is limited because free electrons provide a strong dissipation channel for spin waves, which leads to lifetimes of only a few nanosecondsKorenman and Prange (1972). Such short lifetimes pose a serious challenge for quantum operations with magnons, but they can still be of interest for studies in cryogenic magnonics Li et al. (2019). Working with lifetimes of this order requires strong coupling between systems to ensure sufficiently fast energy transfer.
Figure 1: (a) Temperature-dependent magnetization of hematite. Inset shows the easy-axis crystal structure for $T<T_{\mathrm{M}}$ (taken from Chen et al. (2025)). (b) Sketch of the antiferromagnetic (AFM) spin texture under microwave excitation. The magnetic field is applied along the $x$ direction. Inset shows a SEM image of the coplanar waveguide (CPW); scale bar $500~\mathrm{nm}$ (taken from Chen et al. (2025)). (c) Altermagnet: $d$-wave–type spin-split bands (upper left) and perovskite schematic with sublattice-dependent, anisotropic spin-dependent hoppings (red/blue arrows) (taken from Naka, Motome, and Seo (2025)). (d) Illustration of spin waves in a 2D van der Waals magnet. Magnon spintronics relies on manipulation and control of magnon spin transport from an injector to a detector. The electrical contacts for generating, controlling, and detecting magnons are shown as gold pads, while the van der Waals magnet is depicted by blue and yellow atoms (magnetic and non-magnetic, respectively). The spin is represented by a blue arrow whose orientation varies spatially to indicate the transported magnon (taken from Mañas-Valero et al. (2025)). (e) Atoms and spins in a bilayer of the van der Waals antiferromagnet CrPS4. Red and blue arrows indicate the local magnetic moments of the Cr atoms (turquoise). The interlayer (intralayer) exchange coupling is ferromagnetic (antiferromagnetic) (taken from de Wal et al. (2023)). (f) Lifetime $\tau$ of secondary dipolar–exchange magnons (DESW) at $f_{\mathrm{FMR}}/2$ as a function of temperature $T$ (logarithmic $T$-axis) for three YIG spheres. Illustration shows a schematic of three-magnon splitting — the $k=0$ uniform mode ($f_{\mathrm{FMR}}=3.17~\mathrm{GHz}$) splits into two counter-propagating magnons at $1.59~\mathrm{GHz}$ with $|k|\approx 3~\mathrm{rad}\,\mu\mathrm{m}^{-1}$ in the dipolar-exchange spin-wave regime (taken and adapted from Serha et al. (2025b)). (g) Experimental FMR linewidth $\Delta f$ as a function of the FMR frequency $f_{\mathrm{FMR}}$ for a $100~\mathrm{nm}$ YIG film on a GGG substrate. The half-solid points are measurement points up to $18~\mathrm{GHz}$ for temperatures down to $2~\mathrm{K}$ and up to $10~\mathrm{GHz}$ for temperatures below $2~\mathrm{K}$ (gray dashed lines mark these limits). The straight lines are Gilbert fits performed up to $18~\mathrm{GHz}$ and $10~\mathrm{GHz}$, respectively. Above these values, the linewidth shows non-Gilbert behavior versus $f_{\mathrm{FMR}}$. The linewidth broadening and deviation from Gilbert behavior are due to magnetization of the GGG substrate at low temperatures (inset), which creates an inhomogeneous stray field affecting the internal field of the YIG film (taken and adapted from Serha et al. (2025c)). (h) Crystal structure of diamagnetic YSGAG, an alternative substrate to paramagnetic GGG for YIG. The cubic lattice structure of YSGAG is similar to that of YIG, but with iron ions replaced at the octahedral sites by scandium, gallium, and aluminum, while tetrahedral sites are occupied by gallium and aluminum (taken from Serha et al. (2025d)).
Heusler compounds
Heusler compounds, a class of mostly ternary or quaternary intermetallics with highly tunable electronic and magnetic properties Graf, Felser, and Parkin (2011), offer an improvement over conventional metallic ferromagnets. Among the most promising is CMFS $(\mathrm{Co_{2}Mn_{0.6}Fe_{0.4}Si})$. It exhibits longer spin-wave propagation lengths and lifetimes for dipolar spin waves,Tsymbal and Žutić (2019), yet these lifetimes remain orders of magnitude shorter than of the magnons currently used in quantum magnonic experiments. The finite electrical conductivity maintains a high spin-wave dissipation rate.
Antiferromagnets
Antiferromagnets have emerged as a promising platform for magnonics and quantum magnonics, driven by a series of recent theoretical advances. Foundational work has shown that antiferromagnetic insulators can support superfluid spin transport, enabling coherent long-range spin currents without net magnetization Takei et al. (2014). Their rich spin textures and ultrafast dynamics further highlight the potential of antiferromagnets for spintronic and magnonic devices Gomonay et al. (2018). More recently, the quantum aspects of antiferromagnetic magnons have been explored. A transmon qubit has been proposed as a probe of the quantum characteristics of antiferromagnetic magnonsAzimi-Mousolou et al. (2023), while spin–orbit interactions have been predicted to mediate their hybridization with plasmons, together opening access to entanglement, nonclassical correlations, and new hybrid excitations with enhanced functionalityDyrdał et al. (2023). A representative antiferromagnet of growing interest in magnonics is hematite ($\alpha$-Fe2O3) and shown in Fig. 1 (a). Bulk studies at GHz frequencies report low Gilbert damping (see Tab. 1) Chen et al. (2025); Hamdi, Posva, and Grundler (2023); Wang et al. (2023b). In the easy-axis phase below the Morin transition, all-electrical spectroscopy reveals nearly gapless antiferromagnetic magnons bound to $180^{\circ}$ domain walls that propagate over micrometre distances with a group velocity of about $6\,\mathrm{km/s}$, as illustrated in Fig. 1 (b). The domain-wall mode exhibits a full width at half maximum (FWHM) frequency linewidth $\Delta f_{\mathrm{FWHM}}\approx 1.2\,\mathrm{GHz}$ and a corresponding lifetime $\tau\approx 0.3\,\mathrm{ns}$ Chen et al. (2025). Another canted antiferromagnet with even better damping characteristics is yttrium orthoferrite ($\mathrm{YFeO_{3}}$) Das et al. (2022). It exhibits linewidths below 10 mT at frequencies around 350 GHz and a temperature of 20 K, corresponding to magnon lifetimes slightly above one nanosecond. Although antiferromagnets can reach terahertz frequencies, which push the quantum regime to higher temperatures (e.g. $1\,\mathrm{THz}\approx 48\,\mathrm{K}$), the relatively short lifetimes currently limit the use of hematite, $\mathrm{YFeO_{3}}$ and other antiferromagnets for quantum magnonics.
Altermagnets
Recently, altermagnets have emerged as a distinct class of magnetic materials beyond conventional ferro- and antiferromagnets (see Fig. 1 (c)), characterized by compensated collinear order in real space and non-relativistic alternating spin splitting in momentum space Song et al. (2025b); Gomonay et al. (2024). This unique duality yields terahertz-frequency, field-robust dynamics with negligible stray fields—features reminiscent of antiferromagnets — while also enabling spin-polarized band structures and novel transport phenomena such as the crystal anomalous Hall effect, the spin-splitter effect, and anisotropic magnon spectra. Depending on the orbital symmetry of the momentum-dependent spin splitting, altermagnets are classified into even-parity (d-, g- and i-wave) orders with nodal sign changes (e.g., $d$-wave materials Song et al. (2025b); Zarzuela et al. (2025) like RuO2 Gomonay et al. (2024), MnTeAmin et al. (2024)). Alltogether, this makes altermagnets attractive for quantum magnonics, although strong magnon–magnon interactions can lead to reduced lifetimes, and challenges in thin-film growth together with high damping limits their near-future applicability.
2D van der Waals and organic magnets
Recently, two-dimensional (2D) van der Waals magnets, illustrated in Fig. 1 (d) and (e), have attracted renewed interest in magnonicsKruglyak et al. (2024). A notable example is the van der Waals antiferromagnet CrPS4 (see Tab. 1 and Fig. 1 (e)), which supports long-distance magnon transportde Wal et al. (2023). Likewise, van der Waals magnets such as CrCl3 exhibit standing spin waves. Experiments resolve modes across a thickness of $20\,\mu\mathrm{m}$, demonstrating control of spin-wave modes in 2D systemsKapoor et al. (2020). Propagating spin waves have also been detected in the 2D van der Waals ferromagnet Fe5GeTe2, despite relatively high magnetic dampingSchulz et al. (2023). These results highlight 2D materials as tunable and versatile platforms for future magnonic applications. Organic magnets also offer promise through chemical tunability and potentially low-damping propagation. For example, V(TCNE)x exhibits low with magnon lifetimes up to 40 nsLiu et al. (2020); McCullian et al. (2020). However, for both van der Waals and organic magnets, the material quality must be improved in order to further increase the lifetime of the magnons required for becoming a medium candidate for single-magnon experiments.
Hexaferrites
For long magnon lifetimes and extended free paths, it is essential to use electrically insulating materials to eliminate the dissipation channel from magnons into free electrons. A prominent class of such magnetic insulators are the M-typ hexaferrites, with barium hexaferrite (BaFe12O19, BaM) as their flagship representative. BaM is an insulating ferrimagnet with robust performance across a wide temperature range ($T_{\mathrm{C}}\approx 725$ K) Harris (2012), high coercivity and remanence, and a saturation magnetization of $\mu_{0}M_{\mathrm{s}}\approx 380$ kA/m at RT Harris (2012). Its strong intrinsic magnetic anisotropy ($\mu_{0}H_{\mathrm{a}}\approx 1353$ kA/m) enables broadband frequency operation (50–65 GHz in undoped films at low magnetic fields Song et al. (2010), and 20–100 GHz in doped films Levchenko et al. (2024) under moderate biasing magnetic fields <0.8<0.8 T), as well as comparatively low magnetic losses (Gilbert damping constants as low as $\alpha\sim 7\times 10^{-4}$Li et al. (2016); Malkinski (2012). BaM supports SW with high group velocities, while its perpendicular magnetic anisotropy permits isotropic spin-wave propagation in the forward-volume geometry under low magnetic bias. Such advantages make it a highly attractive material for magnonic radio frequency (RF) and data-processing applications (e.g., for 5G FR25G tools for RF Wireless (2025); Levchenko et al. (2024) 24.25 GHz to 71.0 GHz frequency band), yet magnon lifetimes between several and tens of nanoseconds remain insufficient for experiments that require single-magnon operations. Furthermore, BaM is still mostly limited to macro- and microscale applications due to the lack of a reliable methodology for reproducible growth of high-quality nm-thick films, although ongoing efforts aim to overcome this challenge.
Yttrium Iron Garnet
The longest magnon lifetimes have so far been achieved in YIG, which remains the only material with sufficiently low damping to sustain magnons long enough for single-magnon excitation, measurement, and manipulation. This outstanding material is a ferrimagnetic insulatorCherepanov, Kolokolov, and L’vov (1993); Serga, Chumak, and Hillebrands (2010) has a well understood magnetization behaviorSerha et al. (2025a) and magnetic anisotropyBobkov and Zavislyak (1997). The damping of coherent magnons arises primarily from scattering on crystallographic and surface defectsSpencer, LeCraw, and Clogston (1959); Melkov et al. (2004), from rare-earth impuritiesMihalceanu et al. (2018); Dillon and Nielsen (1959); Spencer, LeCraw, and Linares (1961); Serha et al. (2025b), and from interactions with thermal magnonsPincus, Sparks, and LeCraw (1961); White and Sparks (1963) (spin-spin interaction) and phononsKasuya and LeCraw (1961); Rückriegel et al. (2014); Streib et al. (2019) (spin-lattice interaction). As summarized in Tab. 2, for example at millikelvin temperatures an FMR linewidth of $0.02\,\mathrm{mT}$ corresponds to a magnon lifetime of about $300\,\mathrm{ns}$Serha et al. (2025b), i.e., the minimally required timescale of a few hundred nanoseconds for single-magnon experimentsLachance-Quirion et al. (2020); Xu et al. (2023). Even so, lifetimes of this order are still sufficient to sustain spin-wave coherence for more than 10 µs Makiuchi et al. (2024) and to support magnon propagation over distances exceeding 100 µm using a 110 nm-thin YIG LPE film.Bensmann et al. (2025).
For advanced quantum magnonic operations, longer magnon lifetimes are required. Previous experiments used the Kittel mode, a uniform precession of the magnetization across the whole sample. Its damping is dominated by two-magnon scattering from lattice and surface defects and is therefore largely temperature independent. On cooling to millikelvin temperatures, the lifetime decreases only slightly compared with RTKosen et al. (2019); Serha et al. (2025b). In contrast, short-wavelength magnons with finite wave vector that are localized inside the sample are insensitive to surface quality. For these dipolar-exchange spin waves (DESW) the dominant dissipation is scattering from thermal magnons and phonons. Recently, Serha et al.Serha et al. (2025b) showed that when this thermal bath is depleted the lifetime of short-wavelength magnons increases by more than an order of magnitude and reaches up to 18 µs. These magnons were generated through the intrinsic nonlinearity of the system via three-magnon splitting, with the lifetime set by the power threshold for entering the nonlinear regimeSuhl (1957).
The lifetime of these secondary magnons increases drastically as the temperature decreases from a few kelvin to the millikelvin range as shown in Fig. 1 (f). Multimagnon scattering on thermal magnons and interactions with thermal phonons vanish as the temperature approaches absolute zero. Surface and lattice defects that give rise to two-magnon scattering do not significantly affect DESW, since these bulk modes have micrometer-scale wavelengths. In theory, the lifetime of these magnons should diverge at millikelvin temperatures. In practice, however, it saturates below about 250 mKSerha et al. (2025b). Interestingly, this saturation occurs at different lifetime levels, with crystal contamination with rare and heavy metal ions becoming the limiting factor. Higher-purity samples exhibit longer lifetimes, indicating that values beyond 20 µs — the dephasing time $T_{2}$ of state-of-the-art superconducting qubits Burnett et al. (2019) — are within reach when YIG is grown from specially purified yttrium oxide ($\mathrm{Y_{2}O_{3}}$), which contains most of the rare-earth and heavy-earth impurities responsible for limiting YIG quality at low temperatures.
In total, DESW exhibit lifetimes at millikelvin temperatures that are long enough to transport quantum information and entanglement in micrometer- and submicrometer-thick films, for example between superconducting qubits. This paves the way for new experiments and applications in quantum magnonics, where magnons act as efficient carriers of quantum information.
III YIG films for quantum magnonics
Table 2: Comparison of bulk YIG and thin YIG films relevant to quantum magnonics.
Property bulk YIGSerha et al. (2025b); Maier-Flaig et al. (2017); Klingler et al. (2017) YIG/GGGSerha et al. (2025d) YIG/YSGGYoussef et al. (2025) YIG/YSGAGSerha et al. (2025d)
Saturation magnetization $M_{\mathrm{{}_{S}}}$ @ RT | $\rightarrow$ 0 K (kA/m) 140 | 200 144 | 205 67 | 95 131 | 184
Gilbert damping $\alpha$ @ RT $3\times 10^{-5}$ $4.3\times 10^{-5}$ $1.0\times 10^{-4}$ $4.3\times 10^{-5}$
Inhomogeneous linewidth broadening $\Delta B_{0}$ @ RT (mT) 00.050-0.05 $0.14$ $0.50$ $0.15$
FMR linewidth $\Delta B\,@\approx 8$ GHz @ RT | $\rightarrow$ 0 K (mT) 0.03| 0.020.03\,|\,0.02 0.19| 0.850.19\,|\,0.85 0.50| 0.750.50\,|\,0.75 0.17| 0.250.17\,|\,0.25
Figure 2: FMR FWHM linewidth $\Delta B$ (measure for magnetic damping) as a function of temperature $T$ for an FMR frequency $f_{\mathrm{FMR}}\approx 8\,\mathrm{GHz}$, displayed on a double logarithmic scale. The graph compares measurements obtained in this work with selected results of sputtered and LPE grown YIG films from relevant literature, enabling a direct evaluation of temperature-dependent damping behavior across different samples, their substrates and studies Will-Cole et al. (2023); Guo et al. (2023); Legrand et al. (2025); Serha et al. (2025d) (taken and adapted from Serha et al. (2025d)).
To use magnons as quantum data carriers in quantum magnonics, it is crucial to move beyond standing magnons in spherical resonatorsLachance-Quirion et al. (2019, 2020); Xu et al. (2023); Li et al. (2022); Rani et al. (2025) and establish YIG films as the active medium for propagating single magnons. At RT, the highest bulk like quality YIG films, from micrometer to nanometer scales, have been grown on paramagnetic GGG substrates because of the nearly lattice matching between the two crystals Dubs et al. (2017, 2020); Ding et al. (2020); Heyroth et al. (2019); Cornelissen et al. (2015); Onbasli et al. (2014); Hahn et al. (2014); Askarzadeh and Shokrollahi (2025).
However, quantum magnonics must operate at very low temperatures down to the millikelvin range. Under these conditions, paramagnetic substrates such as GGG become readily magnetized by an externally applied magnetic field Danilov et al. (1989); Serha et al. (2024). The magnetized substrate generates an inhomogeneous stray field that depends on its geometry and imposes a strong gradient in the internal field of the YIG film. For example, a common $(5\times 5\times 0.5)$ mm3 GGG substrate at 1.8 K and an in-plane bias field of 600 mT produces stray fields of up to 60 mT at the edges and about 12 mT at the center opposing the external fieldSerha et al. (2024). This gradient is already broadening the FMR linewidths measured over the sampleSerha et al. (2025c). By selecting an appropriate substrate geometry or by microstructuring the YIG film, the GGG-induced stray-field gradient can be eliminated across the magnetic medium, resulting in a homogeneous internal field in YIGSchmoll et al. (2025b). However, the partially magnetically ordered substrate couples to the YIG spin system Schmoll et al. (2025a), opening extra dissipation channels that broaden the ferromagnetic resonance (FMR) linewidth Serha et al. (2025c, 2024); Guo et al. (2022); Jermain et al. (2017); Mihalceanu et al. (2018); Kosen et al. (2019) and reduce magnon lifetimes of propagating dipolar magnons Schmoll et al. (2025a). The linewidth broadening effect is shown in Fig. 1 (g)Serha et al. (2025c) and Tab. 2 reports a linewidth difference of more than a factor of four between RT and $T\,\rightarrow 0$, even in the absence of the inhomogeneous GGG-induced stray field. Such a decrease in magnon lifetime undermines the advantages that YIG would otherwise offer for quantum-magnonic experiments. MicrostructuringSchmoll et al. (2025b) and operating at low external fieldsKnauer et al. (2023) can mitigate the stray-field gradient to some extent. Nevertheless, eliminating the coupling between the substrate spin system and the YIG film requires the use of diamagnetic substrates. For a comparison of the temperature-dependent damping behavior between YIG films grown on GGG and on diamagnetic alternative substrates, Fig. 2 presents literature results on sputtered and LPE-grown YIG films from different studies, covering various samples and substratesWill-Cole et al. (2023); Guo et al. (2023); Legrand et al. (2025); Serha et al. (2025d).
Yttrium aluminum garnet (YAG) was among the first diamagnetic substrates proposed to replace paramagnetic GGG in YIG heterostructures, aiming to unlock new possibilities for quantum magnonics at cryogenic temperatures by avoiding substrate magnetization in applied fields. However, despite the drawbacks of GGG at low temperatures, YIG/GGG still outperforms YIG/YAG at both RT and low temperature (see Fig. 2), primarily because the large lattice mismatch between YIG and YAG degrades crystalline quality of the YIG filmSposito et al. (2013); Krysztofik et al. (2021); Serha et al. (2025d).
In general, the broad garnet family Enke et al. (1978); Yoshimoto et al. (2019) offers promising diamagnetic substrates with lattice constants closely matching YIG. For example, yttrium scandium aluminum garnet (YSAG), used as a diamagnetic spacer between YIG and GGG, has been shown to reduce damping in YIG films at low temperatures Guo et al. (2022). A stronger candidate is yttrium scandium gallium garnet (YSGG), which can serve directly as a substrateGuo et al. (2023); Legrand et al. (2025). YIG/YSGG has caught up with YIG/GGG in terms of FMR linewidth at low temperaturesYoussef et al. (2025) (see Tab. 2) and even demonstrated robust and superior spin-wave propagation at 2 KAbrão et al. (2025). However, this improvement comes at the cost of lower saturation magnetization, which yields slower dipolar magnons, as the group velocity of magetostatical spin waves are pratically directly proportional to the magnetization saturation. In addition, for YIG/YSGG, only thin films of about 60 nm can be grown fully strained due to the large lattice mismatch, and these still do not reach the best room-temperature damping values achieved by YIG grown on GGG substratesSerha et al. (2025d).
A breakthrough in the field came with the introduction of yttrium scandium gallium aluminum garnet (YSGAG) by C. Guguschev and C. Dubs et al. Guguschev et al. (2025). This diamagnetic garnet enables improved lattice matching up to zero mismatch with YIG. Figure 1 (h) illustrates the $\mathrm{Y}_{3}\mathrm{(Sc,Ga,Al)}_{2}\mathrm{(Ga,Al)}_{3}\mathrm{O}_{12}$ crystal structure, highlighting selected dodecahedral, octahedral, and tetrahedral oxygen coordination coordination polyhedra with the corresponding ions at their centers. By customizing the distribution ratio of the diamagnetic ions Sc, Ga, and Al on the octahedral and tetrahedral sites, the lattice parameter can be tuned to match that of epitaxial YIG films Guguschev et al. (2025). As shown in Fig. 2, taken from the studySerha et al. (2025d), YIG films grown by liquid-phase epitaxy (LPE) on YSGAG exhibit the same low damping at room temperature and at millikelvin temperatures that YIG on GGG achieves only at room temperature, thereby eliminating the drawbacks at cryogenic temperatures associated with paramagnetic substrates. Using YSGAG as the substrate, record-low FMR damping in YIG films were observed also at RT, with FMR linewidths as narrow as $0.1\,\mathrm{mT}$ at frequencies up to $4\,\mathrm{GHz}$ and a Gilbert damping parameter $\alpha=4.29\times 10^{-5}$.
From the properties summarized in Tab. 2, it is clear that the YIG/YSGAG system comes closest to bulk YIG at cryogenic temperatures. Its magnon lifetime is still shorter by a factor of about 5–10 compared with ultra-pure bulk YIG Serha et al. (2025b). However, YIG LPE film quality is known to improve with increasing thicknessesDubs et al. (2017), and careful co-tuning of the YIG and YSGAG lattice parameters can yield an essentially perfect match of both systems. This approach enables YIG films with bulk-like magnetic quality at millikelvin temperatures—an ideal platform for studying single-magnon propagation in quantum magnonics.
IV Outlook
Two converging breakthrough are leading to the perspective using ultra long living propagating magnons for future solid-state quantum technologies. First, diamagnetic YSGAG substrates with a compositionally tunable lattice can be matched to YIG for nanometer-thin as well as micrometer-thick films. This approach removes substrate magnetization and stray-field gradients at cryogenic temperatures while preserving the room-temperature damping performance that has made YIG/GGG films the benchmark. Second, in ultra pure YIG up to 18 µs microsecond lifetimes for short wavelength dipolar–exchange magnons were demonstrated once the thermal magnon bath is depleted, showing that long coherence is available when impurities and extrinsic scattering are suppressed. Together these results point to YIG films that support single magnon excitation, their propagation, and manipulation on chip opening new directions for solid-state quantum magnonics.
A practical route to this goal is provided by nanometer-thin YIG films grown on YSGAG, whose lattice parameters can be compositionally tuned to match YIG. For submicrometer- or micrometer-thick films, an even better lattice match between the substrate and YIG is required. Such a perfectly matched, diamagnetic system would avoid the cryogenic drawbacks of paramagnetic GGG or strained substrate/YIG combinations, allowing the film to approach the intrinsic low dissipation of bulk YIG. In this regime, magnon lifetimes are governed mainly by material imperfections, making growth and processing critical: impurity levels must approach ultra-pure single-crystal quality with strict control of rare-earth contamination, while careful surface and edge preparation suppresses two-magnon scattering. At millikelvin temperatures, depletion of the thermal magnon bath allows dipolar-exchange magnons to reach lifetimes above ten microseconds. A remaining obstacle is that DESW achieve long lifetimes by being insensitive to surfaces, but this also implies the low group velocity of dipolar waves, and their lifetime is expected to decrease with the increase in velocity. One potential solution would be to use low-frequency exchange waves with wavelengths shorter than the DESW, with reduced two-magnon scattering efficiencies and high group velocities.
Taken together, these elements will provide access to long-living and long-propagating single magnons, to be used as magnonic quantum information carriers. These carriers can serve as entanglement busses or gates between superconducting qubits placed on the same YSGAG substrates. Consequently, such single magnons will enable the integration of ’classical magnonics’ data processing approaches, including RF applications, Boolean and neuromorphic computing, with magnon-based quantum computing concepts on the same nanoscale chip.
Acknowledgements.
This research was funded in part by the Austrian Science Fund (FWF) project Paramagnonics [10.55776/I6568]. The work was supported by the German Federal Ministry of Research, Technology and Space (BMFTR) under the reference numbers (13N17109) within a collaborative project “Low-loss materials for integrated magnonic-superconducting quantum technologies (MagSQuant)”. The authors thank Khrystyna O. Levchenko for insightful consultations and valuable scientific discussions on magnetic materials, in particular on hexaferrites.
Data Availability Statement
Data sharing is not applicable to this article as no new data were created or analyzed in this study.